OpenStudy (anonymous):

Find the derivative of f(y)= ln(ln(2y^3))

OpenStudy (anonymous):

I understand its the chain rule, but i'm having difficulty finding the correct answer supplied by the book.

OpenStudy (reemii):

\((\ln(u))' = \frac1{u}u'\). where \(u(y)=\ln(2y^3)\). Now compute \(u'\)... \(u'=(\ln(w))' = \frac1ww'\), where \(w=2y^3\). Now compute \(w'\). and you'll be done.

OpenStudy (anonymous):

So, for the first, its 1/(ln(2y^3)) * ln(2y^3)?

OpenStudy (anonymous):

I suppose my difficulty lies in determining what f(x) and g(x) are in relation to this problem, given that the chain rule is f '(g(x)) * g '(x)

OpenStudy (reemii):

in f(g(x)), f is the "outer" function. it is the \(\ln\) function. g is the argument of that function. it can still be a compound function. you then just use the formula again.

OpenStudy (anonymous):

So, I should end up with 1/(ln(2y^3)) * 1/(2y^3) * 6y^2?

OpenStudy (reemii):

I think so.

OpenStudy (reemii):

and you can make simplifications.

OpenStudy (reemii):

don't forget that \(\ln(x^\alpha) = \alpha\ln(x)\).

OpenStudy (anonymous):

thats what I was missing, thank you very much!

OpenStudy (reemii):

yw